INFRARED OBSERVATION OF THE ν_{2}(σ) STRETCHING MODE OF LINEAR GeC_{3}

\underline{E. GONZALEZ}, C. M. L. RITTBY, and W. R. M. GRAHAM, Department of Physics and Astronomy, Texas Christian University, Fort Worth, TX 76129.

GeC_{3}Ge was first^{1} produced by dual laser ablation of germanium and carbon rods, and the ν_{3}(σ_{u}) stretching fundamental was assigned at 1920.7 cm^{−1}. Later work^{2} enhanced the production of the molecule via laser ablation of a single, sintered germanium-carbon rod, thus enabling the identification of two additional infrared active vibrational fundamentals ν_{4}(σ_{u})=735.6 cm^{−1} and ν_{6}(π_{u})=580.1 cm^{−1}. In the present work using the same technique, GeC_{3} trapped in solid Ar at ∼10 K has been observed by FTIR spectroscopy . Comparison of ^{13}C isotopic shift measurements with the predictions of density functional theory (DFT) calculations at the B3LYP/cc-pV(D,T)Z level confirm the identification of the ν_{2}(σ) stretching fundamental at 1279.6 cm^{−1}.

1 D.L. Robbins, C.M.L. Rittby, and W.R.M. Graham, J. Chem. Phys. 114, 3570 (2001). 2 E. Gonzalez, C.M.L. Rittby, and W.R.M. Graham, J. Chem. Phys. in press.